The duration of the AstraZeneca Covid-19 vaccine side effect is normally distributed with a mean of 1.5 days and a standard deviation of 0.4 days.


(i) If a AstraZeneca Covid-19 vaccine recipient is selected at random, find the percentage

that the duration of side effect is less than 24 hours. (3 marks)


(ii) 96% of the AstraZeneca Covid-19 vaccine recipients have experienced more than k days of discomfort symptoms. Find the value of k. (3 marks)


(iii) 7 AstraZeneca Covid-19 vaccine recipients are selected randomly, find the probability

that 3 recipients have less than 24 hours of side effect.


1
Expert's answer
2021-10-12T15:52:08-0400

i)

"z=\frac{x-\mu}{\sigma}=\frac{1-1.5}{0.4}=-1.25"


"P(x<1)=P(z<-1.25)=0.10565"


ii)

"P(x>k)=0.96"

"P(x<k)=1-0.96=0.04"

z for 0.04:


"z=\frac{k-\mu}{\sigma}=\frac{k-1.5}{0.4}=-1.25"

"k=1.5-1.25\cdot 0.4=1" day


iii)

Binomial distribution

"p=P(x<1)=0.10565" (from part i))

"n=7"

"P(k=3)=C^3_7p^3q^4=\frac{7!}{3!4!}\cdot0.10565^3\cdot(1-0.10565)^4=0.0264"